# Cross-Band Channel Impulse Response Prediction: Leveraging 3.5 GHz Channels for Upper Mid-Band

### v1

@@ I Introduction
Figure 1: Architecture of CIR-UNext: Gain-UNext is an AU-Net-Aux with a $(4\times L+1)$-channel input for efficient and accurate per-path gain prediction, where $L$ is the number of resolved multipath components. Phase-UNext shares the simliar AU-Net backbone (without the red auxiliary part), with a $(2\times L)$-channel input, but is trained separately for per-path phase prediction. In our experiments, we consider $L=20$.
@@ II-A Physical Motivation for Cross-Band Inference
where $L$ is the number of multipath components and $\tau_{l}$ is the propagation delay of the $l$-th path. Each per-path term can be modeled as
@@
where ${\bf f}_{\text{rx}}(\cdot)\!\in\!{\mathbb{C}}^{2\times 1}$ and ${\bf f}_{\text{tx}}(\cdot)\!\in\!{\mathbb{C}}^{2\times 1}$ denote the antenna field patterns at the specified azimuth and elevation angles, accounting for both vertical and horizontal polarizations. The per-path transfer matrix ${\bf T}_{l}(f)\!\in\!{\mathbb{C}}^{2\times 2}$ encapsulates all propagation effects along the $l$-th path. Here, $k=2\pi/\lambda$ is the wavenumber, with $\lambda$ denoting the wavelength; $({\bf r}_{\text{rx},l},{\bf r}_{\text{tx},l})$ are the unit direction vectors of arrival and departure, respectively, and $({\bf d}_{\text{rx},u},{\bf d}_{\text{tx},s})$ represent the relative positions of the receive and transmit antenna elements with respect to their array centers.
@@ II-B Ray-Tracing Dataset Generation
In (3), the amplitude term $\lambda/(4\pi d_{l})$ represents the free-space path loss, while material absorption and scattering efficiency further influence both the magnitude and phase through the interaction matrices $\mathbf{C}_{m}(f)$. The phase factor $e^{-jkd_{l}}$ accounts for the propagation delay, exhibiting a faster phase rotation at higher frequencies for a fixed $d_{l}$. Additional phase shifts are introduced by the reflection and scattering matrices $\mathbf{C}_{m}(f)$.
@@ II-C Deep Learning for Cross-Band Mapping
The dataset is partitioned at the building-map level into training, validation, and testing sets in an 80%–10%–10% ratio. Data augmentation, including rotation and mirroring, is applied to improve robustness and generalization. The models are trained using mean square error (MSE) loss for 200 epochs on an NVIDIA RTX 5080 GPU. Testing is performed on 900 test samples, equally divided between two building scenarios: BM1 (open area) and BM2 (dense urban), as illustrated in Fig. 2. Notably, both BM1 and BM2 are unseen during training.
@@
Figure 2: Building Maps used in the experiments. (a) BM1, characterized by a prominent open space. (b) BM2, featuring densely clustered buildings.
@@ III-A Different Gain-UNext Architectures
To evaluate performance, we employ the average absolute error over the entire area as the evaluation metric, defined as
$\text{G-Err}=\frac{1}{P_{\text{eff}}\times L}\sum_{p=1}^{P_{\text{eff}}}\sum_{l=1}^{L}\left|\hat{x}_{p,l}-x_{p,l}\right|,$ (4)
where $P_{\text{eff}}$ denotes the number of valid positions (excluding building-occupied areas), $L=20$ is the number of predicted paths, $\hat{x}_{p,l}$ is the predicted path gain at the $p$-th position of the $l$-th path (corresponding to $|h_{s,u,l}(f)|$ in (2)), and $x_{p,l}$ is the corresponding ground-truth value. All path gains are expressed in decibels (dB).
@@ III-B Useful Information for Cross-Band Path Gain Prediction
where $G(\phi)$ is the horizontal antenna pattern, $G_{0}$ is the boresight gain, $\phi_{3\text{dB}}$ is the half-power beamwidth, and $\phi$ is the angular offset from the antenna boresight. Since path gain is also strongly correlated with the AoD, including AoD as an additional input feature can further improve prediction performance.
Table II summarizes the results under different input settings with the same standard AU-Net architecture. The Baseline configuration, using only 3.5 GHz path gain and interaction number, yields a median error of $1.04$ dB (IQR $1.99$ dB). Adding antenna pattern information reduces the error to $0.95$ dB (IQR $1.79$ dB). Incorporating both antenna pattern and AoD further improves performance, lowering the median error to $0.70$ dB and the IQR to $1.34$ dB. These results confirm the effectiveness of antenna pattern and directional features when considering different antenna patterns in cross-band path gain prediction.
@@ IV-A Propagation Distance Phase Shift as Input
To evaluate performance, we employ the average absolute phase error over an area as the evaluation metric, defined as
$\text{P-Err}=\frac{1}{P_{\text{eff}}L}\sum_{p=1}^{P_{\text{eff}}}\sum_{l=1}^{L}\min\!\left\{|\hat{\phi}_{p,l}-\phi_{p,l}|,\;2\pi-|\hat{\phi}_{p,l}-\phi_{p,l}|\right\},$ (6)
@@ Acknowledgment
This work was supported in part by the National Science and Technology Council of Taiwan under grant NSTC 114-2221-E-110-031-MY3 and the SNS JU project 6G-DISAC under the EU’s Horizon Europe research and innovation programme under Grant Agreement No 101139130.

### v2

@@ I Introduction
Figure 1: Architecture of CIR-UNext: Gain-UNext is an AU-Net-Aux with a $(4\times L+1)$-channel input for efficient and accurate per-path gain prediction, where $L$ is the number of resolved multipath components. Phase-UNext shares the simliar AU-Net backbone (without the auxiliary part), with a $(2\times L)$-channel input, but is trained separately for per-path phase prediction. In our experiments, we consider $L=20$.
@@ II-A Physical Motivation for Cross-Band Inference
where $L$ is the number of multipath and $\tau_{l}$ is the delay of the $l$-th path. Each per-path term can be modeled as
@@
where ${\bf f}_{\text{rx}}(\cdot)\!\in\!{\mathbb{C}}^{2\times 1}$ and ${\bf f}_{\text{tx}}(\cdot)\!\in\!{\mathbb{C}}^{2\times 1}$ denote the antenna field patterns at the specified azimuth and elevation angles. The per-path transfer matrix ${\bf T}_{l}(f)\!\in\!{\mathbb{C}}^{2\times 2}$ encapsulates all propagation effects along the $l$-th path. Here, $k=2\pi/\lambda$ is the wavenumber, with $\lambda$ denoting the wavelength; $({\bf r}_{\text{rx},l},{\bf r}_{\text{tx},l})$ are the unit direction vectors of arrival and departure, respectively, and $({\bf d}_{\text{rx},u},{\bf d}_{\text{tx},s})$ represent the relative positions of the receive and transmit antenna elements with respect to their array centers.
@@ II-B Ray-Tracing Dataset Generation
In (3), the term $\lambda/(4\pi d_{l})$ represents the free-space path loss, while material absorption and scattering efficiency further influence both the magnitude and phase through the interaction matrices $\mathbf{C}_{m}(f)$. The phase factor $e^{-jkd_{l}}$ accounts for the propagation delay, exhibiting a faster phase rotation at higher frequencies for a fixed $d_{l}$. Additional phase shifts are introduced by the reflection and scattering matrices $\mathbf{C}_{m}(f)$.
@@ II-C Deep Learning for Cross-Band Mapping
The dataset is split at the building-map level into training, validation, and testing sets in an 80%–10%–10% ratio. Rotation and mirroring are used for data augmentation to enhance robustness and generalization. The models are trained using mean square error (MSE) loss for 200 epochs on an NVIDIA RTX 5080 GPU. Testing is performed on 900 test samples, equally divided between two building scenarios: BM1 (open area) and BM2 (dense urban), as illustrated in Fig. 2.
@@
Figure 2: Building Maps used in the experiments. (a) BM1, characterized by a prominent open space. (b) BM2, featuring densely clustered buildings. Notably, both BM1 and BM2 are unseen during training.
@@ III-A Different Gain-UNext Architectures
To evaluate performance, we employ the absolute error over the entire area as the evaluation metric, defined as
$\text{G-Err}=\left|\hat{x}_{p,l}-x_{p,l}\right|,$ (4)
where $p$ denotes the valid positions (excluding building-occupied areas), $\hat{x}_{p,l}$ is the predicted path gain at the $p$-th position of the $l$-th path (corresponding to $|h_{s,u,l}(f)|$ in (2)), and $x_{p,l}$ is the corresponding ground-truth value. All path gains are expressed in decibels (dB).
@@ III-B Useful Information for Cross-Band Path Gain Prediction
where $G_{0}$ is the boresight gain, $\phi_{3\text{dB}}$ is the half-power beamwidth, and $\phi$ is the angular offset from the antenna boresight. Since path gain is also strongly correlated with the AoD, including AoD as an additional input feature can further improve prediction performance.
Table II summarizes the results under different input settings with the same standard AU-Net architecture. The Baseline configuration, using only 3.5 GHz path gain and interaction number, yields a median error of $1.04$ dB (IQR $1.99$ dB). Adding antenna pattern reduces the error to $0.95$ dB (IQR $1.79$ dB). Incorporating both antenna pattern and AoD further improves performance, lowering the median error to $0.70$ dB and the IQR to $1.34$ dB. These results confirm the effectiveness of antenna pattern and directional features when considering different antenna patterns in cross-band path gain prediction.
@@ IV-A Propagation Distance Phase Shift as Input
To evaluate performance, we employ the absolute phase error over an area as the evaluation metric, defined as
$\text{P-Err}=\min\!\left\{|\hat{\phi}_{p,l}-\phi_{p,l}|,\;2\pi-|\hat{\phi}_{p,l}-\phi_{p,l}|\right\},$ (6)
@@ Acknowledgment
This work was supported in part by the National Science and Technology Council (NSTC) of Taiwan under Grants NSTC 114-2221-E-110-031-MY3, NSTC 114-2218-E-110-005, and NSTC 113-2222-E-002-010, and in part by the SNS JU project 6G-DISAC under the EU’s Horizon Europe research and innovation programme under Grant Agreement No. 101139130.